MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2007/06/27
OBSERVER: Fukui Nagaya Paul
CONDITION: clear
COMMENT: We took a special field for MOA268 as often as gb5,9. DARK D2405(60sec), D2406(200sec) are OK. We exchanged the broken vacuum-sensor (cold-cathode type) for the new one (hot-cathode type).
Compressor gas pressure: 1.8 MPa
Expected Observation Time: START 18:22 END 07:19


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB8006-80595432919740
LMC1578-1578161200
SMC392-39212400
ALERT346-3571212720
GRB-000
TOTAL---6834922060 (06h07m40s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
131313135310111154101111
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
111310111010101010110


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
111111000000
Field13Field14Field15Field16
0000


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
110000000000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun
ALERTMOA2007BLG26812


CHECK SHEET

TimeStatusWeather
18:30observingclear, a special field were taken for MOA268
19:00observingclear, a special field were taken for MOA268
20:00observingclear, a special field were taken for MOA268
21:00observingclear, return to normal mode
22:00observingclear
23:00observingclear
0:00observingclear
1:00observingclear
2:00observingslightly cloudy
3:00waitingcloudy
4:00waitingcloudy
5:00waitingcloudy
6:00observingpartly cloudy
7:00observingpartly cloudy


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day